# Some Fibonacci-Related Sequences

### 18 Dec 2023 v1

@@ 3 Finite automata and regular expressions
A finite automaton is a simple model of a computer. In its simplest incarnation, a deterministic finite automaton (DFA), the automaton can be in any one of a finite number of different states. It takes as input a finite string over a finite alphabet, and processes its input symbol-by-symbol, choosing the next state through a lookup table (the transition function) based only on the current state and the next input symbol. Some of the states are designated as accepting states; an input string is accepted if the automaton is in an accepting state after all its symbols are processed. In this paper we represent integers in Fibonacci representation, so we can think of the DFA as processing integers instead of strings. Thus, a DFA is a good model to represent a set of natural numbers; namely, the set of all integers whose representations are accepted. DFA’s are typically drawn as transition diagrams, where the arrows represent transitions between states, double circles denote accepting states, and single circles denote nonaccepting states.
@@ 5 Guessing and checking
One of the principal tools we use is the “guess-and-check” method. Here we guess a finite automaton computing a sequence based on empirical data, and then use Walnut to verify that it is correct, usually by checking that the defining equation holds. The guessing procedure is described in [8, §5.7].
@@
and Walnut returns TRUE. ∎
Proposition 5.
Proof.
We use the following Walnut command:
def a007067 "?msd_fib Ex \$phin(2*n,x) & z=(x+1)/2":
def a007064 "?msd_fib ~Ej \$a007067(j,n)":
eval checkp2 "?msd_fib An (Ek \$p2(k,n)) <=> \$a007064(n)":
@@ 7 Missing numbers
We have
$\lfloor\varphi^{2}n+1/2\rfloor$ $=\lfloor(\varphi^{2}2n+1)/2\rfloor$
$=\left\lfloor{{\lfloor\varphi^{2}2n+1\rfloor}\over 2}\right\rfloor$
$=\left\lfloor{{{\lfloor\varphi^{2}2n\rfloor}+1}\over 2}\right\rfloor,$
where we have used the fact that $\lfloor{{\lfloor x\rfloor}\over n}\rfloor=\lfloor x/n\rfloor$ for real numbers $x$ and integers $n\geq 1$. We can now use Walnut to define a synchronized automaton for A004937:
@@ 9 Sorting the terms
We just sketch the proof, which depends on linear representations. For more about these, see [8, Chap. 9] and [2]. Given automata for $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$, we can find linear representations counting the number of occurrences of each natural number $n$ in each sequence. Then we can form the linear representation for the difference of these two. Then we use the fact that whether a linear representation is identically 00 is decidable. ∎

### 22 Dec 2023 v2

@@ 3 Finite automata and regular expressions
A finite automaton is a simple model of a computer. In its simplest incarnation, a deterministic finite automaton (DFA), the automaton can be in any one of a finite number of different states. It takes as input a finite string over a finite alphabet, and processes its input symbol-by-symbol, choosing the next state through a lookup table (the transition function) based only on the current state and the next input symbol. Some of the states are designated as accepting states; an input string is accepted if the automaton is in an accepting state after all its symbols are processed. In this paper we represent integers in Fibonacci representation, so we can think of the DFA as processing integers instead of strings. Thus, a DFA is a good model to represent a set of natural numbers; namely, the set of all integers whose representations are accepted. In this sense, a DFA decides membership for a set (or sequence). DFA’s are typically drawn as transition diagrams, where the arrows represent transitions between states, double circles denote accepting states, and single circles denote nonaccepting states.
@@
We note in passing that, due to a nice result of Charlier et al. [3], the minimal DFAO computing $(n\bmod k)_{n\geq 0}$ for $n$ in Fibonacci representation is known to have $2k^{2}$ states.
@@
The most powerful of all three representations for a sequence is the synchronized automaton, since one can obtain the other two from it easily. The downside is that it is quite possible that there exists a DFA to decide membership in a sequence $\bf s$, but there is no corresponding synchronized automaton computing the $n$’th term of $\bf s$.
@@ 5 Guessing and checking
One of the principal tools we use is the “guess-and-check” method. Here we guess a finite automaton computing a sequence based on empirical data, and then use Walnut to verify that it is correct, usually by checking that the defining equation holds. The guessing procedure is described in [9, §5.7].
@@
Proof.
The sequence A007067 is defined to be $\lfloor n\varphi+1/2\rfloor$. Now
$\lfloor n\varphi+1/2\rfloor=\left\lfloor{{2n\varphi+1}\over 2}\right\rfloor\\ =\left\lfloor{{\lfloor 2n\varphi+1\rfloor}\over 2}\right\rfloor\\ =\left\lfloor{{\lfloor 2n\varphi\rfloor+1}\over 2}\right\rfloor,$
where we have used the fact that $\lfloor{{\lfloor x\rfloor}\over n}\rfloor=\lfloor x/n\rfloor$ for real numbers $x$ and integers $n\geq 1$. Hence A007067 can be computed by the following Walnut code:
def a007067 "?msd_fib Ex \$phin(2*n,x) & z=(x+1)/2":
The sequence A007064 consists of those integers not in A007067. It is not hard to see that A007064 is given (after a change of index) by $n\rightarrow\lfloor n\varphi+\varphi/2\rfloor+n+1$ for $n\geq 0$. Indeed, we can verify this as follows:
def a007067 "?msd_fib Ex \$phin(2*n,x) & z=(x+1)/2":
def a007064 "?msd_fib Ex \$phin(2*n+1,x) & z=n+1+x/2":
eval check_two "?msd_fib Ax (Em \$a007067(m,x)) <=> (~En \$a007064(n,x))":
Finally we can prove the result about $p_{2}$ as follows:
eval checkp2 "?msd_fib An (Ek \$p2(k,n)) <=> (Em \$a007064(m,n))":
and Walnut returns TRUE. ∎
Proposition 5.
@@ 7 Missing numbers
We have, using the same idea as in the proof of Proposition 4, that
$\lfloor\varphi^{2}n+1/2\rfloor=\left\lfloor{{{\lfloor\varphi^{2}2n\rfloor}+1}\over 2}\right\rfloor.$
We can now use Walnut to define a synchronized automaton for A004937:
@@ 9 Sorting the terms
We just sketch the proof, which depends on linear representations. For more about these, see [9, Chap. 9] and [2]. Given automata for $(a(n))_{n\geq 0}$ and $(b(n))_{n\geq 0}$, we can find linear representations counting the number of occurrences of each natural number $n$ in each sequence. Then we can form the linear representation for the difference of these two. Then we use the fact that whether a linear representation is identically 00 is decidable. ∎
@@ 13 Other results
For the remainder of this section, let $a(n)$ represent (as before) the $n$’th term of A105774, let $b(n)=\lfloor n\varphi\rfloor$, and let $c(n)=\lfloor n\varphi^{2}+1/2\rfloor$. We will find some relations concerning the composition of the functions $a$, $b$, and $c$.
Theorem 23.
  • (a)
    $a(b(n))\geq b(a(n))$ for all $n\geq 0$.
  • (b)
    For $n\geq 1$, the difference $a(b(n))-b(a(n))$ is exactly the parity of the number of $1$’s in the Fibonacci representation of $n-1$, that is, sequence A095076.
  • (c)
    $b(a(b(n)))\geq a(b(a(n)))$ for all $n\geq 0$.
  • (d)
    $a(a(b(n)))-a(b(a(n)))\in\{-2,0,2\}$ for all $n\geq 0$.
  • (e)
    If $d(n):=c(a(n))-a(b(n))-a(n)$, then $d(n)\in\{-1,0,1\}$ for all $n\geq 0$.
  • (f)
    For $n\geq 1$ we have $d(n)=0$ iff ${\bf f}[n-1]=1$.
  • (g)
    For $n\geq 1$ we have $c(a(b(n)))+a(b(n))-a(b(b(n)))-2b(a(n))=1$.
Proof.
We use the following Walnut commands:
reg even1 msd_fib "(0*10*1)*0*":
# even number of 1’s in Fibonacci representation
def ab "?msd_fib Ex \$phin(n,x) & \$a105774(x,z)":
# z = a105774(floor(phi*n))
def ba "?msd_fib Ex \$a105774(n,x) & \$phin(x,z)":
# z = floor(phi*a105774(n))
# check that ab >= ba for all n so subtraction makes sense over N:
def test "?msd_fib An,x,y (\$ab(n,x) & \$ba(n,y)) => x>=y":
# Walnut says TRUE
def abmba "?msd_fib Ex,y \$ab(n,x) & \$ba(n,y) & z=x-y":
# 12 states
#test that a(b(n)) = b(a(n)) precisely when parity of 1’s in
#Fibonacci expansion of n-1 is 0:
eval test1 "?msd_fib An \$abmba(n+1,0) <=> \$even1(n)":
# Walnut says TRUE
#so it follows immediately
# that a(b(n)) > b(a(n)) precisely when parity of 1’s in
#Fibonacci expansion of n-1 is 1, no need to test it
def aba "?msd_fib Ex \$ba(n,x) & \$a105774(x,z)":
# 38 states
def bab "?msd_fib Ex \$ab(n,x) & \$phin(x,z)":
# 18 states
# check the inequality bab(n) >= aba(n):
eval test3 "?msd_fib An,x,y (\$bab(n,x) & \$aba(n,y)) => x>=y":
# Walnut says true
def aab "?msd_fib Ex \$ab(n,x) & \$a105774(x,z)":
# 34 states
# check the proposition that aab(n) - aba(n) takes values in
# {-2,0,2} only:
eval test4 "?msd_fib An,x,y (\$aab(n,x) & \$aba(n,y)) => (x=y|x=y+2|y=x+2)":
# Walnut says TRUE
def a004937 "?msd_fib Ex \$phi2n(2*n,x) & z=(x+1)/2":
def ca "?msd_fib Ex \$a105774(n,x) & \$a004937(x,z)":
def dp "?msd_fib Ew,x,y \$ca(n,w) & \$ab(n,x) & \$a105774(n,y) & z+x+y=w+1":
# here dp computes d(n)+1
eval test1 "?msd_fib An,x \$dp(n,x) => (x=0|x=1|x=2)":
eval test2 "?msd_fib An (n>=1) => (F[n-1]=@1 <=> \$dp(n,1))":
def cab "?msd_fib Ex \$ab(n,x) & \$a004937(x,z)":
def abb "?msd_fib Ex \$phin(n,x) & \$ab(x,z)":
eval test3 "?msd_fib An,r,s,t,u (n>=1 & \$cab(n,r) & \$abb(n,s) & \$ab(n,t)
   & \$ba(n,u)) => r+t=s+2*u+1":